# Perturbation Theory and the Quantum Rabi-model

### v1

@@ 7. The Weyl law for some generalizations of the Quantum Rabi Model
In this section we deal with the asymptotic spectral properties of some generalizations of the Quantum Rabi Model. We extend to this setting the $2$-term asymptotics of the Weyl spectral counting function obtained in Malagutti and Parmeggiani [12] for semiregular SMGES systems (see Definition 3.5) modelled after the Jaynes-Cummings Model. In that paper the Jaynes-Cummings Model is written as a semiregular Non-Commutative Harmonic Oscillator, in the SMGES class as introduced in [12]: studying the high-energy spectral properties of this class of pseudodifferential systems, a Weyl law and a refined Weyl law (Theorems 7.8 and 7.9 in [12]) were obtained. (For the same class, also the meromorphic continuation of the spectral zeta-function [11], and a spectral quasi-clustering property [13]have been obtained.) In the present paper, adopting a similar approach, we show how to obtain a Weyl law for a class of SMGES systems that includes and generalizes the QRM.
@@ 7.1. Quantum Rabi Model generalizations
In this section we introduce a class of semiregular NCHOs which generalizes the Quantum Rabi Model. For these classes we state and prove a Weyl law result in Section 7.2 below. Let $E_{jk}$ be the $N\times N$ matrix which is $1$ in the $jk$ entry and 0 otherwise. In what follows $p_{2}^{\mathrm{w}}(x,D)=(|D|^{2}+|x|^{2})/2$ will be the $n$-dimensional quantum harmonic oscillator. One has the following list of models that generalize the QRM, in the different instances in analogy with the Jaynes-Cummings systems [8].
@@ 7.1.1. The Quantum Rabi Model for an $N$-level atom and n=N1n=N-1 cavity-modes in the $\Xi$-configuration
In this case, for $\alpha>0,$ $\alpha,\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@
In this case, for $\alpha>0,$ $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@
In this case, for $\alpha>0,$ $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@ 7.2. A Weyl law for the generalized Quantum Rabi Models
(n:=N1n:=N-1) where $a_{k}=a_{k}^{*}$ is positively homogeneous of degree $k$. Suppose that $a_{2}=p_{2}I_{N}$, where now $p_{2}(X)=|X|^{2}/2,$ $X\in\mathbb{R}^{2n}$, is the $n$-d scalar harmonic oscillator and that on $\dot{\mathbb{R}^{2n}}$ there is a smooth matrix-valued function $b_{1}$, positively homogeneous of degree $1$, such that $a_{\epsilon}:=a+\varepsilon b_{1}$ is an SMGES for all $\varepsilon\in(0,1)$. Suppose that $A=a^{\mathrm{w}}(x,D)$, realized as an unbounded operator in $L^{2}$ with maximal domain $B^{2}$, is positive.
@@
as $\lambda\rightarrow+\infty$.
@@
To prove the theorem we use a perturbation argument. We consider $A=a^{\mathrm{w}}(x,D)$ realized as an unbounded operator in $L^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$ with maximal domain $D(A)=B^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$, and write $B_{1}=b_{1}^{\mathrm{w}}(x,D)$ realized also as an unbounded operator in $L^{2}$ with domain $B^{1}$. We then let $A_{\varepsilon}=A+\varepsilon B_{1}$ realized on the domain $D(A_{\varepsilon})=D(A)$ (independent of $\varepsilon$). The point will be to obtain an operator inequality between $A_{\varepsilon}$ and $A$ which, by the minimax principle, leads to a spectral inequality between $A_{\varepsilon}$ and $A$. The perturbation $A_{\varepsilon}$ belongs to the class SMGES for $\varepsilon>0$. Next, we use Theorem 7.8 (Weyl law) in [12] to have a Weyl law for $A_{\varepsilon}$. Finally, we can obtain a Weyl law for $A$ by the spectral inequality just proven. Actually, the Weyl law for the second order operator $A$ is not refined since the perturbation has order $1$, whence only the first term after the leading term of the asymptotics can be determined precisely.
@@
as $\lambda\rightarrow+\infty$, where constant in $O_{\varepsilon}(\lambda)$ may depend on $\varepsilon$. In a similar way, for $\mathsf{N}_{A_{\varepsilon}}\circ\nu_{\varepsilon,-}^{-1}$ we have
@@
$\leq{\displaystyle\limsup_{\lambda\rightarrow+\infty}}\left(\frac{1}{\lambda^{n-1/2}}\left(\mathsf{N}_{A}(\lambda)-(2\pi)^{-n}\lambda^{n}N\int_{A^{1/2}\leq 1}dX\right)\right)$

### v2

@@ 7. The Weyl law for some generalizations of the Quantum Rabi Model
In this section we deal with the asymptotic spectral properties of some generalizations of the Quantum Rabi Model. We extend to this setting the $2$-term asymptotics of the Weyl spectral counting function obtained in Malagutti and Parmeggiani [13] for semiregular SMGES systems (see Definition 3.5 above) modelled after the Jaynes-Cummings Model. In that paper the Jaynes-Cummings Model is written as a semiregular Non-Commutative Harmonic Oscillator, in the SMGES class as introduced in [13]: studying the high-energy spectral properties of this class of pseudodifferential systems, a Weyl law and a refined Weyl law (Theorems 7.8 and 7.9 in [13]) were obtained. (For the same class, also the meromorphic continuation of the spectral zeta-function [12], and a spectral quasi-clustering property [14] have been obtained.) We next show how to obtain a Weyl law for a class of systems that includes and generalizes the QRM.
@@ 7.1. Quantum Rabi Model generalizations
In this section we introduce a class of semiregular NCHOs which generalizes the Quantum Rabi Model. For this class we state and prove a Weyl law result in Section 7.2 below. Let $E_{jk}$ be the $N\times N$ matrix which is $1$ in the $jk$ entry and 0 otherwise. In what follows $p_{2}^{\mathrm{w}}(x,D)=(|D|^{2}+|x|^{2})/2$ will be the $n$-dimensional quantum harmonic oscillator. One has the following list of models that generalize the QRM, in the different instances in analogy with the Jaynes-Cummings systems [8].
@@ 7.1.1. The Quantum Rabi Model for an $N$-level atom and n=N1n=N-1 cavity-modes in the $\Xi$-configuration
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@
In this case, for $\alpha_{1},\ldots\alpha_{N-1}\in\mathbb{R}\setminus\{0\}$, $\gamma_{1},\ldots\gamma_{N-1}\in\mathbb{R}$ with $\gamma_{1}\leq\gamma_{2}\leq\ldots\leq\gamma_{N-1}$, we consider the $N\times N$ system in $\mathbb{R}^{n}$, n=N1n=N-1, given by
@@ 7.2. A Weyl law for the generalized Quantum Rabi Models
(n=N1n=N-1) where $a_{k}=a_{k}^{*}$ is positively homogeneous of degree $k$. Suppose that $a_{2}=p_{2}I_{N}$, where now $p_{2}(X)=|X|^{2}/2,$ $X\in\mathbb{R}^{2n}$, is the $n$-d scalar harmonic oscillator and that on $\dot{\mathbb{R}}^{2n}$ there is a smooth matrix-valued function $b_{1}$, positively homogeneous of degree $1$, such that $a_{\epsilon}:=a+\varepsilon b_{1}$ is an SMGES for all $\varepsilon\in(0,1)$. Suppose that $A=a^{\mathrm{w}}(x,D)$, realized as an unbounded operator in $L^{2}$ with maximal domain $B^{2}$, is positive.
@@
as $\lambda\rightarrow+\infty$. (Here $ds$ is the Riemannian measure induced on $\{p_{2}=1\}$ by the $2n$-dimensional Euclidean one.)
@@
To prove the theorem we use a perturbation argument. We consider $A=a^{\mathrm{w}}(x,D)$ realized as an unbounded operator in $L^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$ with maximal domain $D(A)=B^{2}(\mathbb{R}^{n};\mathbb{C}^{N})$, and write $B_{1}=b_{1}^{\mathrm{w}}(x,D)$ realized also as an unbounded operator in $L^{2}$ with domain $B^{1}(\mathbb{R}^{n};\mathbb{C}^{N})$. We then let $A_{\varepsilon}=A+\varepsilon B_{1}$ realized on the domain $D(A_{\varepsilon})=D(A)$ (independent of $\varepsilon$). The point will be to obtain an operator inequality between $A_{\varepsilon}$ and $A$ which, by the minimax principle, leads to a spectral inequality between $A_{\varepsilon}$ and $A$. The perturbation $A_{\varepsilon}$ belongs to the class SMGES for $\varepsilon>0$. Next, we use Theorem 7.8 (Weyl law) in [13] to have a Weyl law for $A_{\varepsilon}$. Finally, we can obtain a Weyl law for $A$ by the spectral inequality just proven. Actually, the Weyl law for the second order operator $A$ is not refined since the perturbation has order $1$, whence only the first term after the leading term of the asymptotics can be determined precisely.
@@
as $\lambda\rightarrow+\infty$, where the constant in $O_{\varepsilon}(\lambda)$ may depend on $\varepsilon$. In a similar way, for $\mathsf{N}_{A_{\varepsilon}}\circ\nu_{\varepsilon,-}^{-1}$ we have